Assume: Demand Curve: QD = 80 – 10P; and Supply Curve: QS = 10P 7. Given the information derived above, identify on the graph consumer surplus and producer surplus for each situation as well as deadweight loss, if any. What is the total surplus? Quantify the quantity of goods imported. describe the implications of the imposition of a tariff in this market F. now assume a tariff of $1.00 is added to the world price of $2.00. INCLUDE A GRAPH
Answer #1


Demand. QD = 80 - 10P Supply Qs = 10P At equilibrium, QD=Qs 80 - 10P = 10P 20P = 80 P= 80/20 = $4 Q = 10P = 10*4 = 40 units EConsumer surplus CS = 1 * 60* (8-2) Consumer surplus CS = 1/2*60*6 = 180 Producer surplus PS = 1/2*20* 2 = 20 Total surplus =price (8) world price 20 40 50 60 CS - Consumer Surplus PS - Producer Surplus R - tax revenue D - dead weight loss